# The stable trees revisited

### v1

@@ 1.1 The $\alpha$-stable trees
The $\alpha$-stable trees are a family of random $\mathbb{R}$-trees which constitute the scaling limits of the size-conditioned family trees of Bienaymé branching processes whose offspring distribution is critical and lies in the domain of attraction of an $\alpha$-stable law, for $\alpha\in(1,2]$. The $2$-stable tree is a (constant multiple of) the Brownian continuum random tree (see Aldous [aldous1, aldous2, aldous3]); the rest of the family was introduced by Duquesne and Le Gall [DuquesneLeGall, DuquesneLeGallProbFract], building on earlier work of Le Gall and Le Jan [LeGallLeJan]. The focus of this paper is the case $\alpha\in(1,2)$.
@@
$\mathbb{E}\left[\exp(-\lambda L_{t})\right]=\exp(-t\lambda^{\alpha}),\quad\lambda\geq 0,t\geq 0.$
@@
and an equivalence relation by declaring $x\sim y$ if $d^{\circ}_{\alpha}(x,y)=0$. Finally, take $\mathcal{T}_{\alpha}$ to be the quotient space $0,1]/{\sim}$, which is endowed with the metric $d_{\alpha}$ induced on $\mathcal{T}_{\alpha}$ by $d_{\alpha}^{\circ}$. Take $\mu_{\alpha}$ to be the push-forward of the Lebesgue measure on $[0,1]$ onto $\mathcal{T}_{\alpha}$. (See [LeGall, DuquesneLeGall, duquesnealphalimit] for more details.)
@@
Let $T^{n}$ be the rooted ordered tree which represents the genealogy of a branching process with offspring distributed as $\xi$, conditioned to have size $n$. We think of this as a metric space by endowing its vertices $V(T^{n})$ with the graph distance, $d^{n}$, and additionally we let $\mu^{n}$ be the uniform probability measure on its vertices. Theorem 3.1 of Duquesne [duquesnealphalimit] proves the convergence of the height and contour processes of $T^{n}$ to their continuous analogues (see also Kortchemski [igorduquesneproof] for a streamlined proof). The following result is a corollary.
@@
Haas and Miermont [haasmiermbt] provide an alternative approach to the proof of Theorem 1.1 in the case where $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ for some constant $c>0$ (see their Theorem 8).
@@ 1.2 Line-breaking construction
$Z_{k}=\inf\{t>0:\tau_{t}>U_{i}\tau_{C_{k}-}\}.$
(This picks out the time of a jump of $\tau$ in $(0,Y_{k})$ chosen with probability proportional to its size. Note that, almost surely, $Y_{k}$ is not a jump-time of the process $\tau$, and so $\tau_{Y_{k}}=\tau_{Y_{k}-}$ a.s.) If the cut and attachment points are generated in this way, we say that $(\mathcal{T},d)$ is generated by a line-breaking construction with intensity process $\tau$. The case $\tau_{t}=t$, where $Z_{k}$ is uniformly distributed on $0,Y_{k}]$, is precisely Aldous’ line-breaking construction of the Brownian CRT [aldous1].
@@
Along the way, we show that $(\mathcal{T}(k))_{k\geq 1}$ realises the random finite-dimensional distributions of the $\alpha$-stable tree. That is, for each $k\geq 1$, $\mathcal{T}(k)$ has the same distribution as the subtree of $\mathcal{T}_{\alpha}$ spanned by the root and $k$ vertices chosen independently according to the measure $\mu_{\alpha}$. (As shown by Aldous [aldous3], the random finite-dimensional distributions characterise the law of any continuum random tree.)
@@ 1.3 Discussion
Line-breaking constructions of random $\mathbb{R}$-trees date right back to Aldous’ first paper on the Brownian continuum random tree [aldous1]. That construction was generalised by Aldous and Pitman [aldouspitmanicrt] to give the family of inhomogeneous continuum random trees (ICRTs), which were further studied in [camarripitmanicrt, aldouspitmanicrt2, aldousicrtexplore, abrcompactness, abrfennec]. This family is parametrised by a sequence $\theta=(\theta_{i})_{i\geq 0}$ such that $\theta_{1}\geq\theta_{2}\geq\dots\geq 0$, $\sum_{i\geq 0}\theta_{i}^{2}=1$ and either $\theta_{0}\neq 0$ or $\sum_{i=1}\theta_{i}=\infty$. For $i\geq 1$, let $E_{i}\sim\mathrm{Exp}(\theta_{i})$ and then set
@@
Then use $(\tau_{t})_{t\geq 0}$ as the intensity process in a line-breaking construction. (This succinct formulation, which is due to Blanc-Renaudie [abrcompactness], is equivalent to the original formulation of [aldouspitmanicrt].) The ICRTs were shown by Camarri and Pitman [camarripitmanicrt] to be the scaling limits of a family of discrete models called $p$-trees. We will find it helpful in the sequel to relax the condition $\sum_{i\geq 0}\theta_{i}^{2}=1$ to the condition $\|\theta\|_{2}<\infty$, the effect being simply a scaling of all edge-lengths in the tree by $\|\theta\|_{2}$. Let us write $(\mathcal{T}^{\theta},d^{\theta},\mu^{\theta})$ for the ICRT with parameter $\theta$. Blanc-Renaudie [abrcompactness] proves that $(\mathcal{T}^{\theta},d^{\theta})$ is a.s. compact if and only if
@@
Other line-breaking constructions with general lengths and uniform attachment mechanism have been studied in [AminiDevroyeGriffithsOlver, CurienHaas, Haas].
A first, more complicated, line-breaking construction for the $\alpha$-stable trees was given by the first author with Bénédicte Haas in [lbcstable] in 2015. This construction is not based on an inhomogeneous Poisson point process, but rather on an increasing Markov chain, $(M_{k})_{k\geq 1}$. In order to describe it, we need first to introduce the two-parameter family of generalised Mittag-Leffler distributions, $\mathrm{ML}(\theta,\beta)$, for $\beta\in(0,1)$ and $\theta>-\beta$. If $X\sim\mathrm{ML}(\theta,\beta)$ then $X$ takes values in $\mathbb{R}^{+}$ and has distribution determined by its moments,
@@
(A second version of the construction was also given in [lbcstable] but since it does not apparently reveal any extra information about the stable tree, we do not describe it here. See also Rembart and Winkel [RembartWinkel] for an alternative perspective.)
In [lbcstable], it is proved that $(\widetilde{\mathcal{T}}(k))_{k\geq 1}$ realises the random finite-dimensional distributions of the stable tree. Therefore, the construction makes manifest the distributions of various distances in the tree (see Proposition 1.6 of [lbcstable] for more details). We are able to recover these facts using our new construction, although it requires some work, for example:
@@
The use of line-breaking constructions has recently acquired new impetus, beginning with work of Addario-Berry, Blanc-Renaudie, Donderwinkel, Maazoun and Martin [surveypaper], who gave a survey of analogous discrete constructions for uniform random trees with a given degree sequence. We describe a particular instance of this in detail in Section 3 below, and with randomised degrees it will play a key role in our proof of Theorem 1.1. (It has also recently been used to study various aspects of random trees with given degrees in [abrfixeddegrees, ABDrandomtrees, snakes, ABDKcriticaltrees].)
It is possible to consider mixtures of ICRTs by randomising the parameter $\theta$: that is, we first sample $\theta$ according to some distribution and then, conditionally given $\theta$, we sample an ICRT with parameter $\theta$. (In this context, we do not wish to constrain $\|\theta\|_{2}$ to take the value 1.) Aldous, Pitman and Miermont [aldousicrtexplore] conjectured that the Lévy trees of Duquesne and Le Gall [DuquesneLeGall] (which include the stable trees) are mixtures of ICRTs. This statement has been proved in the case of the stable trees by Wang [minmin]. (We understand that a complete proof in the general case may appear in a future version of the preprint [abrfixeddegrees].)
Theorem 1.4 (Wang [minmin]).
@@ 2.1 A measure-changed subordinator
Let $a,b\geq 0$, and let $\nu$ be a measure on $\mathbb{R}^{+}$ such that $\int_{\mathbb{R}^{+}}(x\wedge x^{2})\,\nu(dx)<\infty$. Let $L=(L_{t})_{t\geq 0}$ be a spectrally positive Lévy process in $\mathbb{R}$ with characteristic exponent
@@
We will only illustrate the case $a=b=0$ as it reduces notation (but exactly the same argument still works in the general case — we just have more terms to keep track of). Condition (ii) implies that $e^{-\Psi_{L}(\lambda)}$ is in $L^{1}$, so by Fourier inversion $L$ has a density, namely
@@
In what follows we write $M_{t}=M_{t}^{\nu,a,b}$ for brevity. Let $(\mathcal{F}_{t})$ be the filtration associated to $\sigma$. We will reduce the martingale property to an equivalent statement about (unconditional) expectations. Let $t,s\geq 0$ and define $\sigma_{r}^{\prime}=\sigma_{s+r}-\sigma_{s}$; then $\sigma^{\prime}$ has the same law as $\sigma$ and is independent of $\mathcal{F}_{s}$, and we may compute
@@
We can now handle the expectation using Campbell’s formula: indeed, $\sigma$ is a pure jump subordinator, so we can find a Poisson process on $\mathbb{R}^{+}\times\mathbb{R}^{+}$, of intensity $ds\otimes\nu(dx)$, such that the set of atoms $\Pi\subset\mathbb{R}^{+}\times\mathbb{R}^{+}$ satisfies
@@
$\mathbb{E}\left[\exp\left(\int_{0}^{t}(c+\sigma_{r})\,dr\right)p(-c-\sigma_{t})\right]\!$ $=\frac{1}{2\pi}\int_{\mathbb{R}}\mathbb{E}\left[\exp\left(i\lambda\sigma_{t}+\int_{0}^{t}\sigma_{r}\,dr\right)\right]e^{c(t+i\lambda)-\Psi_{L}(\lambda)}\,d\lambda$
$=\frac{1}{2\pi}\int_{\mathbb{R}}\!\exp\Big(\Psi_{L}(\lambda)-\Psi_{L}(\lambda-it)+c(t+i\lambda)-\Psi_{L}(\lambda)\Big)\,d\lambda$
$=\frac{1}{2\pi}\int_{\mathbb{R}}\exp\Big(ci(\lambda-it)-\Psi_{L}(\lambda-it)\Big)\,d\lambda$
$=\frac{1}{2\pi}\int_{\mathbb{R}}\exp(G(\lambda-it))\,d\lambda.$
@@
For each $\epsilon\in(0,1]$ define a measure $\nu_{\varepsilon}$ on $(0,\infty)$ by $\nu_{\varepsilon}(dx)=C_{\alpha}x^{-\alpha-1}e^{-\varepsilon x^{2}}\,dx$. We apply Theorem 2.1 in the case $a=b=0,\nu=\nu_{\varepsilon}$: let $L^{\varepsilon}$ and $\sigma^{\varepsilon}$ denote the relevant Lévy process and subordinator.
@@
By monotone convergence, the integral in the above expression converges to a finite quantity $I_{\alpha}=\int_{0}^{\infty}(1-\cos(u))u^{-\alpha-1}$ as $|\lambda|\to\infty$, and so $\textnormal{Re}\left(\Psi_{L^{\varepsilon}}(\lambda)\right)\sim CI_{\alpha}|\lambda|^{\alpha}$ as $|\lambda|\to\infty$. In particular, $\textnormal{Re}\left(\Psi_{L^{\varepsilon}}\right)\gg|\lambda|$.
Hence, letting $p_{\varepsilon}$ be the density of $L_{1}^{\varepsilon}$, we see that $M^{\varepsilon}_{t}=\exp(\int_{0}^{t}\sigma_{s}^{\varepsilon})p_{\varepsilon}(-\sigma_{t}^{\varepsilon})$ is a martingale with mean $p_{\varepsilon}(0)$. As seen in the sketch proof of Theorem 2.1, it is easier to work with the condition
@@
and deduce the analogous statement for $\sigma$ by taking limits. In what follows, we consider $c\geq 0$ fixed. To handle the expectation on the left-hand side, it is helpful to consider a coupling of $\sigma^{\varepsilon}$ and $\sigma$. Since $\nu_{\varepsilon}\ll\nu$ with Radon–Nikodym derivative $e^{-\varepsilon x^{2}}$ taking values in $[0,1]$, we can achieve this by thinning: consider a Poisson process $\Pi\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\times[0,1]$ with atoms denoted by $(x,t,u)$ and intensity measure $\nu\otimes dt\otimes du$. Then we set
@@
$|p_{\varepsilon}(x)-p(x)|$ $=\left|\frac{1}{2\pi}\int_{\mathbb{R}}e^{-i\lambda x}\left(\exp(-\Psi_{L_{\varepsilon}}(\lambda))-\exp(-\Psi_{L}(\lambda))\right)\,d\lambda\right|$
$\leq\frac{1}{2\pi}\int_{\mathbb{R}}\left|e^{-\Psi_{L_{\varepsilon}}(\lambda)}-e^{-\Psi_{L}(\lambda)}\right|\,d\lambda,$
a bound which does not depend on $x$. The integrand tends to 0 as $\varepsilon\downarrow 0$ and, by our calculations above, is bounded by $2e^{-\delta|\lambda|^{\alpha}\mathbbm{1}\left(|\lambda|\geq 1\right)}$ for some $\delta>0$, so by dominated convergence this integral tends to 0. Hence, we may bound
@@
Now by Theorem 14.35 of Sato [satobook], there exist constants $C_{1},C_{2}>0$ and $\beta=\frac{2-\alpha}{2\alpha-2}$ such that
@@
$\exp\left(\int_{0}^{t}(a+\sigma_{s}^{\varepsilon})\,ds\right)p_{\varepsilon}(-a-\sigma_{t}^{\varepsilon})$ $\leq\exp\left(t(a+\sigma_{t}^{\varepsilon})\right)p_{\varepsilon}(-a-\sigma_{t}^{\varepsilon})$
$\leq M\exp\left(t(a+\sigma_{t}^{\varepsilon})\right)p(-a-\sigma_{t}^{\varepsilon}-a_{\varepsilon})$

### v2

@@ 1.1 The $\alpha$-stable trees
The $\alpha$-stable trees are a family of random $\mathbb{R}$-trees which constitute the scaling limits of the size-conditioned family trees of Bienaymé branching processes whose offspring distribution is critical and lies in the domain of attraction of an $\alpha$-stable law, for $\alpha\in(1,2]$. The $2$-stable tree is a (constant multiple of) the Brownian continuum random tree (see Aldous [9, 10, 11]); the rest of the family was introduced by Duquesne and Le Gall [22, 23], building on earlier work of Le Gall and Le Jan [34]. The focus of this paper is the case $\alpha\in(1,2)$.
@@
$\mathbb{E}\left[\exp(-\lambda L_{t})\right]=\exp(t\lambda^{\alpha}),\quad\lambda\geq 0,t\geq 0.$
@@
and an equivalence relation by declaring $x\sim y$ if $d^{\circ}_{\alpha}(x,y)=0$. Finally, take $\mathcal{T}_{\alpha}$ to be the quotient space $0,1]/{\sim}$, which is endowed with the metric $d_{\alpha}$ induced on $\mathcal{T}_{\alpha}$ by $d_{\alpha}^{\circ}$. Take $\mu_{\alpha}$ to be the push-forward of the Lebesgue measure on $[0,1]$ onto $\mathcal{T}_{\alpha}$. (See [35, 22, 24] for more details.)
@@
Let $T^{n}$ be the rooted ordered tree which represents the genealogy of a branching process with offspring distributed as $\xi$, conditioned to have size $n$. We think of this as a metric space by endowing its vertices $V(T^{n})$ with the graph distance, $d^{n}$, and additionally we let $\mu^{n}$ be the uniform probability measure on its vertices. Theorem 3.1 of Duquesne [24] proves the convergence of the height and contour processes of $T^{n}$ to their continuous analogues (see also Kortchemski [33] for a streamlined proof). The following result is a corollary.
@@
Haas and Miermont [28] provide an alternative approach to the proof of Theorem 1.1 in the case where $\mathbb{P}\left(\xi=k\right)\sim ck^{-\alpha-1}$ for some constant $c>0$ (see their Theorem 8).
@@ 1.2 Line-breaking construction
$Z_{k}=\inf\{t>0:\tau_{t}>U_{k}\tau_{Y_{k}-}\}.$
(This picks out the time of a jump of $\tau$ in $(0,Y_{k})$ chosen with probability proportional to its size. Note that, almost surely, $Y_{k}$ is not a jump-time of the process $\tau$, and so $\tau_{Y_{k}}=\tau_{Y_{k}-}$ a.s.) If the cut and attachment points are generated in this way, we say that $(\mathcal{T},d)$ is generated by a line-breaking construction with intensity process $\tau$. The case $\tau_{t}=t$, where $Z_{k}$ is uniformly distributed on $0,Y_{k}]$, is precisely Aldous’ line-breaking construction of the Brownian CRT [9].
@@
Along the way, we show that $(\mathcal{T}(k))_{k\geq 1}$ realises the random finite-dimensional distributions of the $\alpha$-stable tree. That is, for each $k\geq 1$, $\mathcal{T}(k)$ has the same distribution as the subtree of $\mathcal{T}_{\alpha}$ spanned by the root and $k$ vertices chosen independently according to the measure $\mu_{\alpha}$. (As shown by Aldous [11], the random finite-dimensional distributions characterise the law of any continuum random tree.)
@@ 1.3 Discussion
Line-breaking constructions of random $\mathbb{R}$-trees date right back to Aldous’ first paper on the Brownian continuum random tree [9]. That construction was generalised by Aldous and Pitman [8] to give the family of inhomogeneous continuum random trees (ICRTs), which were further studied in [19, 7, 6, 18, 17]. This family is parametrised by a sequence $\theta=(\theta_{i})_{i\geq 0}$ such that $\theta_{1}\geq\theta_{2}\geq\dots\geq 0$, $\sum_{i\geq 0}\theta_{i}^{2}=1$ and either $\theta_{0}\neq 0$ or $\sum_{i=1}\theta_{i}=\infty$. For $i\geq 1$, let $E_{i}\sim\mathrm{Exp}(\theta_{i})$ and then set
@@
Then use $(\tau_{t})_{t\geq 0}$ as the intensity process in a line-breaking construction. (This succinct formulation, which is due to Blanc-Renaudie [18], is equivalent to the original formulation of [8].) The ICRTs were shown by Camarri and Pitman [19] to be the scaling limits of a family of discrete models called $p$-trees. We will find it helpful in the sequel to relax the condition $\sum_{i\geq 0}\theta_{i}^{2}=1$ to the condition $\|\theta\|_{2}<\infty$, the effect being simply a scaling of all edge-lengths in the tree by $\|\theta\|_{2}$. Let us write $(\mathcal{T}^{\theta},d^{\theta},\mu^{\theta})$ for the ICRT with parameter $\theta$. Blanc-Renaudie [18] proves that $(\mathcal{T}^{\theta},d^{\theta})$ is a.s. compact if and only if
@@
Other line-breaking constructions with general lengths and uniform attachment mechanism have been studied in [13, 20, 30].
A first, more complicated, line-breaking construction for the $\alpha$-stable trees was given by the first author with Bénédicte Haas in [27] in 2015. This construction is not based on an inhomogeneous Poisson point process, but rather on an increasing Markov chain, $(M_{k})_{k\geq 1}$. In order to describe it, we need first to introduce the two-parameter family of generalised Mittag-Leffler distributions, $\mathrm{ML}(\theta,\beta)$, for $\beta\in(0,1)$ and $\theta>-\beta$. If $X\sim\mathrm{ML}(\theta,\beta)$ then $X$ takes values in $\mathbb{R}^{+}$ and has distribution determined by its moments,
@@
(A second version of the construction was also given in [27] but since it does not apparently reveal any extra information about the stable tree, we do not describe it here. See also Rembart and Winkel [38] for an alternative perspective.)
In [27], it is proved that $(\widetilde{\mathcal{T}}(k))_{k\geq 1}$ realises the random finite-dimensional distributions of the stable tree. Therefore, the construction makes manifest the distributions of various distances in the tree (see Proposition 1.6 of [27] for more details). We are able to recover these facts using our new construction, although it requires some work, for example:
@@
The use of line-breaking constructions has recently acquired new impetus, beginning with work of Addario-Berry, Blanc-Renaudie, Donderwinkel, Maazoun and Martin [1], who gave a survey of analogous discrete constructions for uniform random trees with a given degree sequence. We describe a particular instance of this in detail in Section 3 below, and with randomised degrees it will play a key role in our proof of Theorem 1.1. (It has also recently been used to study various aspects of random trees with given degrees in [16, 4, 2, 3].)
It is possible to consider mixtures of ICRTs by randomising the parameter $\theta$: that is, we first sample $\theta$ according to some distribution and then, conditionally given $\theta$, we sample an ICRT with parameter $\theta$. (In this context, we do not wish to constrain $\|\theta\|_{2}$ to take the value 1.) Aldous, Pitman and Miermont [6] conjectured that the Lévy trees of Duquesne and Le Gall [22] (which include the stable trees) are mixtures of ICRTs. This has recently been shown by Blanc-Renaudie in [16] for Lévy trees which are Gromov–Prokhorov limits of Bienaymé trees. The corresponding statement in the case of the stable trees was proved by Wang [41] in the following more refined form, which identifies $\theta$ as the sequence of jumps of an $\alpha$-stable excursion.
Theorem 1.4 (Wang [41]).
@@ 2.1 A measure-changed subordinator
Let $a,b\geq 0$, and let $\nu$ be a measure on $\mathbb{R}^{+}$ such that $\int_{\mathbb{R}^{+}}(x\wedge x^{2})\,\nu(dx)<\infty$. Let $L=(L_{t})_{t\geq 0}$ be a spectrally positive Lévy process on $\mathbb{R}$ with characteristic exponent
@@
We will only illustrate the case $a=b=0$ as it reduces notation (but exactly the same argument still works in the general case — we just have more terms to keep track of). Condition (ii) implies that $e^{-\Psi_{L}(\lambda)}$ is in $L^{1}$, so by Fourier inversion $L_{1}$ has a density, namely
@@
In what follows we write $M_{t}=M_{t}^{\nu,a,b}$ for brevity. Let $(\mathcal{F}_{t})$ be the filtration associated to $\sigma$. We will reduce the martingale property to an equivalent statement about (unconditional) expectations. Let $t,s\geq 0$ and define $\sigma_{r}^{\prime}=\sigma_{s+r}-\sigma_{s}$ for $r\geq 0$: then the process $\sigma^{\prime}$ has the same law as $\sigma$ and is independent of $\mathcal{F}_{s}$, and we may compute
@@
We can now handle the expectation using Campbell’s formula: indeed, $\sigma$ is a pure jump subordinator, so we can find a Poisson process on $\mathbb{R}^{+}\times\mathbb{R}^{+}$, of intensity $ds\otimes x\nu(dx)$, for which the set of atoms $\Pi\subset\mathbb{R}^{+}\times\mathbb{R}^{+}$ satisfies
@@
$\mathbb{E}\left[\exp\left(\int_{0}^{t}(c+\sigma_{r})\,dr\right)p(-c-\sigma_{t})\right]$
$\qquad=\frac{1}{2\pi}\int_{\mathbb{R}}\mathbb{E}\left[\exp\left(i\lambda\sigma_{t}+\int_{0}^{t}\sigma_{r}\,dr\right)\right]e^{c(t+i\lambda)-\Psi_{L}(\lambda)}\,d\lambda$
$\qquad=\frac{1}{2\pi}\int_{\mathbb{R}}\!\exp\Big(\Psi_{L}(\lambda)-\Psi_{L}(\lambda-it)+c(t+i\lambda)-\Psi_{L}(\lambda)\Big)\,d\lambda$
$\qquad=\frac{1}{2\pi}\int_{\mathbb{R}}\exp\Big(ic(\lambda-it)-\Psi_{L}(\lambda-it)\Big)\,d\lambda$
$\qquad=\frac{1}{2\pi}\int_{\mathbb{R}}\exp(G(\lambda-it))\,d\lambda.$
@@
For each $\varepsilon\in(0,1]$ define a measure $\nu_{\varepsilon}$ on $(0,\infty)$ by $\nu_{\varepsilon}(dx)=C_{\alpha}x^{-\alpha-1}e^{-\varepsilon x^{2}}\,dx$. We apply Theorem 2.1 in the case $a=b=0,\nu=\nu_{\varepsilon}$: let $L^{\varepsilon}$ and $\sigma^{\varepsilon}$ denote the relevant Lévy process and subordinator.
@@
By monotone convergence, the integral in the above expression converges to a finite quantity $I_{\alpha}=\int_{0}^{\infty}(1-\cos(u))u^{-\alpha-1}\,du$ as $|\lambda|\to\infty$, and so $\textnormal{Re}\left(\Psi_{L^{\varepsilon}}(\lambda)\right)\sim C_{\alpha}I_{\alpha}|\lambda|^{\alpha}$ as $|\lambda|\to\infty$. In particular, $\textnormal{Re}\left(\Psi_{L^{\varepsilon}}\right)\gg|\lambda|$.
Hence, letting $p_{\varepsilon}$ be the density of $L_{1}^{\varepsilon}$, we see that $M^{\varepsilon}_{t}=\exp(\int_{0}^{t}\sigma_{s}^{\varepsilon}\,ds)p_{\varepsilon}(-\sigma_{t}^{\varepsilon})$ is a martingale with mean $p_{\varepsilon}(0)$. As seen in the sketch proof of Theorem 2.1, it is easier to work with the condition
@@
and deduce the analogous statement for $\sigma$ by taking limits. In what follows, we consider $c\geq 0$ fixed. To handle the expectation on the left-hand side, it is helpful to consider a coupling of $\sigma^{\varepsilon}$ and $\sigma$. Since $\nu_{\varepsilon}\ll\nu$ with Radon–Nikodym derivative $e^{-\varepsilon x^{2}}$ taking values in $[0,1]$, we can achieve this by thinning: consider a Poisson point process $\Pi\subset\mathbb{R}^{+}\times\mathbb{R}^{+}\times[0,1]$ with intensity measure $\nu(dx)\otimes dt\otimes du$ and atoms denoted by $(x,t,u)$. Then we set
@@
$|p_{\varepsilon}(x)-p(x)|$ $=\left|\frac{1}{2\pi}\int_{\mathbb{R}}e^{-i\lambda x}\left(\exp(-\Psi_{L^{\varepsilon}}(\lambda))-\exp(-\Psi_{L}(\lambda))\right)\,d\lambda\right|$
$\leq\frac{1}{2\pi}\int_{\mathbb{R}}\left|e^{-\Psi_{L^{\varepsilon}}(\lambda)}-e^{-\Psi_{L}(\lambda)}\right|\,d\lambda,$
a bound which does not depend on $x$. The integrand tends to 0 as $\varepsilon\downarrow 0$ and, by our calculations above, is bounded by $2e^{-\delta|\lambda|^{\alpha}\mathbbm{1}\left(|\lambda|\geq 1\right)}$ for some $\delta>0$, so by dominated convergence this integral tends to 0. Now we may bound
@@
Now by Theorem 14.35 of Sato [39], there exist constants $C_{1},C_{2}>0$ and $\beta=\frac{2-\alpha}{2\alpha-2}$ such that
@@
$\exp\left(\int_{0}^{t}(c+\sigma_{s}^{\varepsilon})\,ds\right)p_{\varepsilon}(-c-\sigma_{t}^{\varepsilon})$ $\leq\exp\left(t(c+\sigma_{t}^{\varepsilon})\right)p_{\varepsilon}(-c-\sigma_{t}^{\varepsilon})$
$\leq M\exp\left(t(c+\sigma_{t}^{\varepsilon})\right)p(-c-\sigma_{t}^{\varepsilon}-a_{\varepsilon})$